OpenStudy (anonymous):

A tangent-tangent angle intercepts two arcs that measure 135 and 225. What is the measure of the tangent-tangent angle?

OpenStudy (arabpride):

i believe its 45 degrees

OpenStudy (ranga):



OpenStudy (ranga):

Angle x = (far arc - near arc) / 2 = (arc ACB - arc ADB) / 2 = (225 - 135) / 2 = 90 / 2 = 45 degrees.

OpenStudy (arabpride):

good explanation^^^

OpenStudy (ranga):

thank you ^^^